OpenStudy (anonymous):

On a highway curve with radius 44 m, the maximum force of static friction (centripetal force) that can act on a 1,023-kg car going around the curve is 7,927 N. What speed limit should be posted for the curve so that cars can negotiate it safely?

OpenStudy (anonymous):

F = m a F = m (v^2/r) 7927 N = (1023) (v^2/44m) gives you max v in m/s, then get km per hour from 1000m = km 3600s = 1 hr.